# Towards Full-Scenario Safety Evaluation of Automated Vehicles: A Volume-Based Method

### v1

@@ 1 Introduction
Automated vehicle (AV) technology has advanced rapidly over the past decade. During this period, an increasing number of commercially available vehicles have been equipped with self-driving capabilities (Li 2022). Although most AVs currently on the market remain at Level 2 automation, several leading AV manufacturers, such as Tesla, Waymo, Huawei, Xpeng, and Li Auto, have recently announced the successful implementation of the ”parking-to-parking” feature. This function enables fully autonomous driving from a parking space at the origin to a parking space at the destination (Huawei 2025). However, the deployment of these technologies in real-world environments presents significant safety challenges (Ding et al. 2023, Zhang et al. 2022). Rigorous and comprehensive safety testing is essential to ensure the reliability, robustness, and public acceptance of both current and future AVs (Moody et al. 2020).
Existing literature on AV safety evaluation primarily focuses on relatively simple and specific driving scenarios. For example, Zhao et al. (2017) and Ma et al. (2023) evaluated the adaptive cruise control function in the car-following scenario. Zhao et al. (2016), Feng et al. (2020b, 2021), and Zhou et al. (2025) extended the testing scenarios to include multiple lanes and lane-changing behaviors. Arvin et al. (2020), Tang et al. (2021) and Song et al. (2022) examined AV behavior at intersections. Feng et al. (2020a) conducted a series of case studies to validate the general framework proposed in Feng et al. (2020c), covering car-following, cut-in, and highway exit scenarios. In addition, Feng et al. (2023) tested a proposed scenario generation method in a roundabout setting. The primary reason for the literature to focus on relatively simple scenarios is that commercial AVs over the past several years have mostly remained at low levels of automation, leading to less urgent demand for evaluation methods targeting complex environments. Furthermore, the lack of high-level AVs has resulted in a shortage of relevant field test data necessary for the development and calibration of AV behavior models. However, with the emergence of high-level AVs, existing evaluation methods designed and tested on car-following and lane-changing scenarios are no longer sufficient to meet the growing safety testing needs.
@@ 4.2 Performance of the Volume Estimation Algorithms in the Single-lane Scenario
We first evaluate the performance of the four algorithms in a single-lane scenario. Since the dimensionality of the problem increases with the time horizon $T$, we test the algorithms under various dimensions, with $T\in\{1,2,3,4,5,10,15,20,25\}$. Here, $T=25$ corresponds to a trajectory length of 5 seconds, which is consistent with the typical scenario duration in the literature. In this section, we use MC to denote the Monte Carlo method. In the results table, columns Percent report the proportion of dangerous scenarios, columns Time report the computation time of each algorithm in seconds, and columns Error report the percentage error of each algorithms’ dangerous scenario proportion compared with the exact results of the VE algorithm. Here we set $\eta=1$ s, which means that the dangerous scenarios have $\mathrm{TTC}\leq 1$. A maximum runtime limit of 3,600 seconds was set on all algorithms. For algorithms that exceed the predefined time limit, the results are marked with \.

### v2

@@ 1 Introduction
Automated vehicle (AV) technology has advanced rapidly over the past decade. During this period, an increasing number of commercially available vehicles have been equipped with self-driving capabilities (Li 2022). Although most AVs currently on the market remain at Level 2 automation, several leading AV manufacturers, such as Tesla, Waymo, Huawei, Xpeng, and Li Auto, have recently announced the successful implementation of the ”parking-to-parking” feature. This feature enables fully autonomous driving from a parking space at the origin to a parking space at the destination (Huawei 2025). However, the deployment of these features in real-world environments presents significant safety challenges (Ding et al. 2023, coppola2023assessing). Rigorous and comprehensive safety testing is essential to ensure the reliability, robustness, and public acceptance of both current and future AVs (Moody et al. 2020, ma2024evolving).
Existing literature on AV safety evaluation primarily focuses on relatively simple and specific driving scenarios. For example, Zhao et al. (2017) and Ma et al. (2023) evaluated the adaptive cruise control function in the car-following scenario. Zhao et al. (2016), Feng et al. (2020b, 2021), and Zhou et al. (2025) extended the testing scenarios to include multiple lanes and lane-changing behaviors. Arvin et al. (2020), Tang et al. (2021) and Song et al. (2022) examined AV behavior at intersections. Feng et al. (2020a) conducted a series of case studies to validate the general framework proposed in Feng et al. (2020c), covering car-following, cut-in, and highway exit scenarios. In addition, Feng et al. (2023) tested a proposed scenario generation method in a roundabout setting. The primary reason for the literature to focus on relatively simple scenarios is that commercial AVs over the past several years have mostly remained at low levels of automation. The limited availability of AVs with high levels of automation has resulted in less urgent demand for evaluation methods targeting complex environments and a shortage of relevant field test data necessary for the development and calibration of AV behavior models. However, as high-level AVs begin to emerge, existing evaluation methods designed and tested on car-following and lane-changing scenarios are no longer sufficient to meet the growing safety testing needs.
@@ 4.2 Performance of the Volume Estimation Algorithms in the Single-lane Scenario
We first evaluate the performance of the four algorithms in a single-lane scenario. Since the dimensionality of the problem increases with the time horizon $T$, we test the algorithms under various dimensions with $T\in\{1,2,3,4,5,10,15,20,25\}$. Here, $T=25$ corresponds to a trajectory length of 5 seconds, which is consistent with the typical scenario duration in the literature. In this section, we use MC to denote the Monte Carlo method. In the results table, columns Percent report the proportion of dangerous scenarios, columns Time report the computation time of each algorithm in seconds, and columns Error report the percentage error of each algorithms’ dangerous scenario proportion compared with the exact results of the VE algorithm. Here we set $\eta=1$ s, which means that the dangerous scenarios have $\mathrm{TTC}\leq 1$. A maximum runtime limit of 3,600 seconds was set on all algorithms. For algorithms that exceed the predefined time limit, the results are marked with \.
@@ 5 Conclusion
6 Acknowledgement
This research was sponsored by the National Science Foundation, CPS: Small: NSF-DST: Turning “Tragedy of the Commons (ToC)” into “Emergent Cooperative Behavior (ECB)” for Automated Vehicles at Intersections with Meta-Learning (No. 2343167).